In a certain federal prison, it is known that 2/3 of the inmates are under 25 years of age. It is also known that 3/5 of the inmates are male and that 5/8 of the inmates are female or 25 years of age or older. What is the probability that a prisoner selected at random from this prison is female and at least 25 years old?


1
Expert's answer
2021-06-17T18:52:40-0400

Let

U- under 25

M- male

F-female

"P(U)=\frac{2}{3}"

"P(M)=\frac{3}{5}"

"P(F\cup\overline U)=\frac{5}{8}"

"P(F\cap\overline U)"

Let

"P(M\cap\overline U)=x"

Then

"P(M\cap{U})=P(M)-P(M\cap \overline U)=\frac{3}{5}-x"

Similarly

"P(F\cap U)=P(U)-P(M\cap U)=\frac{2}{3}-\frac{3}{5}+x=\frac{1}{15}+x"

"P(F\cap\overline U)= P(F)-P(F\cap U)"

"=\frac{5}{8}-x-(\frac{1}{15}-x-)=\frac{67}{120}-2x"

Now

"P(M)+P(F)=1"

"\frac{3}{5}+\frac{5}{8}-x=1"

Solving for x, we have,

"x=\frac{9}{40}"

Thus,

"P(F\cap\overline U)=\frac{67}{120}-2(\frac{9}{40})=\frac{13}{120}"